# paper=a4paper,margin=2.0cm

### v1

@@ Appendix B Asymptotic Constants appearing in Section 4
$\Lambda_{1}=[\sigma_{1}|\alpha_{2}^{(1)}|^{2}(k_{1}^{2}-k_{2}^{2})+\sigma_{2}|\alpha_{2}^{(2)}|^{2}(k_{1}^{2}-k_{2}^{*2})],~{}\hat{\lambda}=(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2},$
$\Lambda_{2}=[\sigma_{1}(k_{1}^{2}-k_{2}^{2})|k_{2}+l_{1}^{*}|^{2}|k_{2}-l_{1}^{*}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}(k_{1}^{2}-k_{2}^{*2})|k_{2}+l_{1}|^{2}|k_{2}-l_{1}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\Lambda_{3}=[\sigma_{1}|k_{1}-k_{2}|^{2}|k_{1}+k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}|k_{2}-l_{1}|^{2}|k_{2}+l_{1}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\Lambda_{4}=[\sigma_{1}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{2}-l_{1}|^{2}|k_{2}+l_{1}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\Lambda_{5}=[\sigma_{1}|k_{1}-k_{2}|^{2}|k_{1}+k_{2}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\Lambda_{6}=[\sigma_{1}(k_{2}^{*2}-l_{1}^{2})|\alpha_{2}^{(1)}|^{2}+\sigma_{2}(k_{2}^{2}-l_{1}^{2})|\alpha_{2}^{(2)}|^{2}],$
$\Lambda_{7}=[\sigma_{1}(k_{2}^{*2}-l_{1}^{2})|k_{1}-k_{2}|^{2}|k_{1}+k_{2}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}(k_{2}^{2}-l_{1}^{2})|\alpha_{2}^{(2)}|^{2}].$ (27)

### v2

@@ Appendix B Asymptotic Constants appearing in Section 4
$\hskip-25.6073pt\Lambda_{1}=[\sigma_{1}|\alpha_{2}^{(1)}|^{2}(k_{1}^{2}-k_{2}^{2})+\sigma_{2}|\alpha_{2}^{(2)}|^{2}(k_{1}^{2}-k_{2}^{*2})],~\hat{\lambda}=(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2},$
$\hskip-25.6073pt\Lambda_{2}=[\sigma_{1}(k_{1}^{2}-k_{2}^{2})|k_{2}+l_{1}^{*}|^{2}|k_{2}-l_{1}^{*}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}(k_{1}^{2}-k_{2}^{*2})|k_{2}+l_{1}|^{2}|k_{2}-l_{1}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\hskip-25.6073pt\Lambda_{3}=[\sigma_{1}|k_{1}-k_{2}|^{2}|k_{1}+k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}|k_{2}-l_{1}|^{2}|k_{2}+l_{1}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\hskip-25.6073pt\Lambda_{4}=[\sigma_{1}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{2}-l_{1}|^{2}|k_{2}+l_{1}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\hskip-25.6073pt\Lambda_{5}=[\sigma_{1}|k_{1}-k_{2}|^{2}|k_{1}+k_{2}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}|\alpha_{2}^{(2)}|^{2}],$
$\hskip-25.6073pt\Lambda_{6}=[\sigma_{1}(k_{2}^{*2}-l_{1}^{2})|\alpha_{2}^{(1)}|^{2}+\sigma_{2}(k_{2}^{2}-l_{1}^{2})|\alpha_{2}^{(2)}|^{2}],$
$\hskip-25.6073pt\Lambda_{7}=[\sigma_{1}(k_{2}^{*2}-l_{1}^{2})|k_{1}-k_{2}|^{2}|k_{1}+k_{2}|^{2}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}(k_{2}^{2}-l_{1}^{2})|\alpha_{2}^{(2)}|^{2}].$ (27)